OpenStudy (anonymous):

Find the rate of change for 2000 to 2010, given these ordered pairs in the data set: (2000, 286) and (2010, 192). Write your answer as a decimal.

OpenStudy (anonymous):

(192-286)/286 = ? over 10 years. Divide by 10 to get rate of change per year.

OpenStudy (ash2326):

Rate of change =\[\frac{192-286}{2010-2000}=\frac{-94}{10}=-9.4\]

OpenStudy (anonymous):

Ash - that is not the "rate" of change. That is the "absolute" value of the average change.

OpenStudy (anonymous):

You need to do the percentage change first and then average it by 10 years.